Jack and Jill Corporation's year-end 20xx balance sheet lists current assets of $268,000, fixed assets of $730,000, current liabilities of $238,000, and long-term debt of $406,000. What is Jack and Jill's total stockholders' equity? (Answer in whole $s, with no $ sign or commas, e.g. xxxxxx, or 123000)
Answer #1

Total assets=current assets+fixed assets

=268,000+730,000=998000

Total liabilities=current liabilities+long term debt

=238,000+406,000=644000

Total assets=Total liabilities+total stockholders' equity

total stockholders' equity=998000-644000

=354000